OpenStudy (anonymous):

Hello everyone can someone help me with this pre cal question Solve the equation by completing the square. 4x2 − 8x + 1 = 0

OpenStudy (chaise):

4x^2 − 8x + 1 = 0 Factor our a 4 4(x^2-2x+1)-3=0 <-- to balance the equation you must minus three to get back down to one after multiplying by four. 4(x-1)^2-3=0 It's hard for me to do this without commentating on what I am doing at the same time. I hope you followed these step. Post another and I will try help.

OpenStudy (anonymous):

where did you get the 3 from? Chaise

OpenStudy (anonymous):

4x2 − 8x + 1 = 0 ===> x2 − 2x + 1/4 = 0 ===> x^2-2x+1=-1/4 + 1 ===> (x-1)^2 = 3/4 ===> x-1 = sqrt 3/2 ===> x=(sqrt3/2 ) +1 or x=(-sqrt3/2) +1

OpenStudy (chaise):

When you expand the bracket you get: 4(x^2-2x+1)-3=0 4x^2-8x+4-3=0 You must -3 so the equation balances. Remember the original was: 4x^2 − 8x + 1 = 0 I hope this made sense?

OpenStudy (chaise):

4(x-1)^2-3=0 Now you can just solve this like a normal equation. Plus three both sides, divide by 4. etcetc

OpenStudy (anonymous):

ok thank you so much for your help i got it. I love this place are there any other sites similar to this or is this the only one?